OpenStudy (anonymous):

Sketch the curve represented by the parametric equations (indicate the direction of the curve). Use a graphing utility to confirm your result. Then eliminate the parameter and write the corresponding rectangular equation whose graph represents the curve x=3-2t y=2+3t

OpenStudy (dumbcow):

by substitution: t = (3-x)/2 then y = 2+3(3-x)/2 which simplifies to y = (13-3x)/2 it is a linear line with slope of -3/2

OpenStudy (anonymous):

so how would my graph look like? and is that what they mean by eliminating the parameter?

OpenStudy (dumbcow):

yes what i did was eliminate the t parameter

OpenStudy (dumbcow):

the graph would be the line

OpenStudy (anonymous):

ok so now i have this question,

OpenStudy (anonymous):

when i do x=sqrt(t) y=1-t be equal to x^2 =t y= 1-x^2

OpenStudy (dumbcow):

correct, though you may have to specify the domain as only positive values

OpenStudy (anonymous):

okk and my graph wouldbe

OpenStudy (dumbcow):

yep

OpenStudy (anonymous):

ok thank you!! :D

OpenStudy (dumbcow):

your welcome